# Concept-Based Interpretable Reinforcement Learning with Limited to No Human Labels

### v1

@@ 5 Related Work
Interpretable Reinforcement Learning Interpretable RL has gained significant attention in recent years [10]. One prominent approach uses rule-based methods — such as decision trees [26, 29], logic [8], and programs [30, 19] — to represent policies. These works either assume that the state is already interpretable or that the policy is pre-specified. Unlike prior work, our method involves learning the interpretable representation (through concept training) for policy learning.
Concept Learning for Reinforcement Learning Inspired by successes in the supervised setting [6, 25, 34], concept-based explanations have recently been incorporated into RL. One approach [7] learns a joint embedding model between state-action pairs and concept-based explanations to expedite learning via reward shaping. Unlike our work, their policy is not a strict function of the concepts, allowing our techniques to be combined to provide both concept-based explanations and a concept-based interpretable policy. Another example, CPM [33], is a multi-agent RL concept architecture that focuses on trade-offs between interpretability and accuracy. They assume that concept labels are available continuously during training. As we have shown, this approach uses over $1$M concept labels in our test environments, whereas our approach reduces the need for continuous human intervention, requiring only 500 or fewer concept labels to achieve similar or better performance in single-agent environments.
Learning Concepts with Human Feedback Another line of work explores how to best leverage human concept labels but not in the RL setting, and does not focus on reducing the labeling burden. In contrast, our approach aims to reduce the concept labeling burden during training. One work [14] instead has users label additional information about the relevance of certain feature dimensions to the concept label. Another work [4] develops an intervention policy at prediction time to choose which concepts to request a label for with the goal of improving the final prediction. Future work could explore using these techniques alongside our method.
@@ 6 Discussion and Conclusion
In this work, we proposed LICORICE, a novel RL algorithm that addresses the critical issue of model interpretability while minimizing the reliance on continuous human annotation. We introduced a training scheme that enables RL algorithms to learn concepts more efficiently from little to no labeled concept data. Our approach interleaves concept learning and RL training, uses an ensemble-based active learning technique to select informative data points for labeling, and uses a simple sampling strategy to better decorrelate the concept data. We demonstrated how this approach reduces manual labeling effort. Finally, we conducted initial experiments to demonstrate how we can leverage powerful VLMs to infer concepts from raw visual inputs without explicit labels in some environments. There are broader societal impacts of our work that must be considered. These include both the impacts of using VLMs in real-world applications, as well as considerations around interpretability more generally. For a more detailed discussion, please refer to Appendix C.
Limitations and Future Work Although VLMs can be successfully used for automatic labeling of some concepts, there are still hallucination issues [2] and other failure cases, such as providing inaccurate counts. We believe that future work that seeks to improve general VLM capabilities and mitigate hallucinations would also help overcome this limitation. Another exciting direction is the work in human-AI complementarity and learning-to-defer algorithms [18] to train an additional classifier for deferring concept labeling to a person when the chance of making an error is high.
Acknowledgments and Disclosure of Funding
This work was supported in part by NSF grant IIS-2046640 (CAREER). Co-author Fang is supported in part by Sloan Research Fellowship. We thank NVIDIA for providing computing resources. We thank Zhicheng Zhang for some helpful discussions, Naveen Raman for pointing us to relevant literature, and Rex Chen and Zhiyu (Zoey) Chen for their constructive comments on a draft of this work.
@@ Appendix A Experimental Result Reproducibility
In this section, we provided detailed descriptions to achieve reproducibility.
A.1 Concepts Definitions
@@
PixelCartPole-v1 Cart Position Continuous $(-2.4,2.4)$ $0.01$\pm$0.0$$
@@
DoorKey-7x7 Agent Position x Discrete 5 $0.36$\pm$0.08$$
@@
DynamicObstacles-5x5 Agent Position x Discrete 3 $0.03$\pm$0.06$$
@@
Table 5: Concepts and their possible values for all environments. For discrete concepts, we report the number of categories. We also provide the mean GPT-4o labeling error (MSELoss for continuous values and 1 - accuracy for discrete ones) for each single concept, averaged across 5 seeds and corresponding to the same evaluation protocols as the ones with largest budgets in Table 8. The $\pm$ [value] part shows the standard deviation. The average errors over concepts are $0.24$\pm$0.06$$, $0.3$\pm$0.02$$, and $0.13$\pm$0.03$$ respectively for the three environments.
Table 5 provides more details on the concepts used in each environment, categorizing them by their names, types, and value ranges. For the PixelCartPole-v1 environment, all concepts such as Cart Position, Cart Velocity, Pole Angle, and Pole Angular Velocity are continuous. In contrast, the DoorKey-7x7 environment features discrete concepts like Agent Position (x and y), Key Position (x and y), and Door Open status, each with specific value ranges. Similarly, the DynamicObstacles-5x5 environment lists discrete concepts, including Agent and Obstacle positions, with corresponding value ranges.
Figure 2: Example start configurations of the DoorKey-7x7 environment. This shows that concepts must general enough to apply to many different configurations. As a result, the agent cannot memorize the exact position of the door and key.
We also visualize the start configurations for DoorKey-7x7 in Figure 2 to illustrate the importance of concept definitions. As shown, the concepts must be defined in such a way to allow the agent to generalize to all possible environment configurations.
A.2 LICORICE Details
@@
Value-Based Methods
@@
Feature Extractor If we use the actor-critic paradigm, we propose to share a feature extractor between the policy and value networks, shown in Figure 3. Intuitively, this choice can offer several advantages compared with using image or predicted concepts as input for both networks.
Sharing a feature extractor enables both networks to benefit from a common, rich representation of the input data, reducing the number of parameters to be trained. More importantly, it balances the updates of the policy and value networks. In experiments, we observed that directly using the raw image as input for both networks complicated policy learning. Conversely, relying solely on predicted concepts for the value network may limit its accuracy in value estimation, particularly if the concepts do not capture all the nuances relevant to the value predictions.
@@ A.3 VLM Details
    Agent Direction: Specify the direction the red triangle is facing, which is the orientation of the vertex (pointy corner) of the isosceles triangle. Choose from ’right’, ’down’, ’left’, or ’up’. Clarify that this direction is independent of movement options.
@@
    • Status: Assess whether the door is open or closed (closed means the door is visible as a whole yellow cell, while open means the door disappears visually). Report as ’true’ for open and ’false’ for closed. Double-check the door’s appearance to confirm if it is open or closed.
@@
    Each direction’s feasibility should be reported as ’true’ if clear and within the grid, and ’false’ otherwise.
@@
    Agent Direction: Specify the direction the red triangle is facing, which is the orientation of the vertex (pointy corner) of the isosceles triangle. Choose from ’right’, ’down’, ’left’, or ’up’. Clarify that this direction is independent of movement options.
@@
    Each direction’s feasibility should be reported as ’true’ if clear and within the grid, and ’false’ otherwise.
@@ A.4 Experimental Details
For the PPO hyperparameters, we set $4\cdot 10^{6}$ total timesteps for PixelCartPole-v1and DoorKey-7x7, and $10^{6}$ for DynamicObstacles-5x5. Besides that, for all environments, we use $8$ vectorized environments, horizon $T=4096$, $10$ epochs for training, batch size of $512$, learning rate $3\cdot 10^{-4}$, entropy coefficient $0.01$, and value function coefficient $1$. For all other hyperparameters, we use the default values from Stable Baselines 3 [23].
For the concept training, we set $100$ epochs with Adam optimizer with the learning rate linearly decaying from $3\cdot 10^{-4}$ to 00 for each iteration in PixelCartPole-v1. In DoorKey-7x7 and DynamicObstacles-5x5, we use the same optimizer and initial learning rate, yet set $50$ epochs instead and set early stopping with threshold linearly increasing from $10$ to $20$, to incentivize the concept network not to overfit in earlier iterations. The batch size is $32$.
For LICORICE, we set the sample acceptance rate $p=0.05$, the ratio for active learning $\tau=10$, batch size to query labels in the active learning module $b=20$, and the number of ensemble models $N=5$. The number of iterations chosen in our algorithm is $M=4$ for PixelCartPole-v1and $M=2$ for the other two environments.
We use NVIDIA A6000 and NVIDIA RTX 6000 Ada Generation. Each of our programs uses less than $2$GB GPU memory. For PixelCartPole-v1 and DoorKey-7x7, each run takes less than 9 hours to finish. For DynamicObstacles-5x5, each run takes less than 2 hours to finish.
For all experiments, we consistently use 5 seeds [$123$, $456$, $789$, $1011$, $1213$] to train the models. We then evaluate on the environment with seed $42$ with 100 episodes and take the average of the reward across the episodes and the concept error across the observations within each episode.
@@ B.1 Balancing Concept Performance and Environment Reward
PixelCartPole-v1 Sequential-Q 0.232 472±0.085 242plus-or-minus0.2324720.085242\$0.232\,472$\pm$\$0.085\,242\$0.232 472 ± 0.085 242 0.102 57±0.037 255plus-or-minus0.102570.037255\$0.102\,57$\pm$\$0.037\,255\$0.102 57 ± 0.037 255
@@
DoorKey-7x7 Sequential-Q 0.462 566±0.091 999plus-or-minus0.4625660.091999\$0.462\,566$\pm$\$0.091\,999\$0.462 566 ± 0.091 999 0.462 327±0.045 955plus-or-minus0.4623270.045955\$0.462\,327$\pm$\$0.045\,955\$0.462 327 ± 0.045 955
@@
DynamicObstacles-5x5 Sequential-Q 0.788 348±0.440 724plus-or-minus0.7883480.440724\$0.788\,348$\pm$\$0.440\,724\$0.788 348 ± 0.440 724 0.014 136±0.009 627plus-or-minus0.0141360.009627\$0.014\,136$\pm$\$0.009\,627\$0.014 136 ± 0.009 627
@@
Table 6: Evaluation of the reward $R$ and concept error achieved by all methods in all environments. This is a extended table from Table 1. The reward is reported as the fraction of the reward upper bound. For PixelCartPole-v1, the $c$ error is the MSE. For the other two environments, the $c$ error is 1 - accuracy. The first four algorithms are given a budget of $B=[500,300,300]$ for each environment, from top to bottom; CPM is given an unlimited budget (in practice, it uses 4M, 4M, 1M concept labels respectively). The $\pm$ [value] part shows the standard deviation. This shows a more complete version of the results in Table 1.
In Table 6, we present an extension of the results in Table 1, including standard deviation. Our method enjoys low variance across all environments in terms of both concept error and reward.

### v2

@@ 5 Related Work
Interpretable RL. Interpretable RL has gained significant attention in recent years (Glanois et al., 2024). One prominent area uses rule-based methods—such as decision trees (Silva et al., 2020; Topin et al., 2021), logic (Delfosse et al., 2024), and programs (Verma et al., 2018; Penkov & Ramamoorthy, 2019)—to represent policies. These works either assume that the state is already interpretable or that the policy is pre-specified. Unlike prior work, our method involves learning the interpretable representation (through concept training) for policy learning.
Concept Learning for RL. Due to successes in the supervised setting (Collins et al., 2023; Sheth & Ebrahimi Kahou, 2023; Zarlenga et al., 2023), concept models have recently been used in RL. Das et al. (2023) trains a joint embedding model between state-action pairs and concept-based explanations to expedite learning via reward shaping. Unlike LICORICE, their policy is not a strict function of the concepts, allowing our techniques to be combined to provide concept-based explanations alongside an interpretable policy. Another example, CPM (Zabounidis et al., 2023), is a multi-agent RL approach that assumes that concept labels are continuously available during training. As we have shown, this approach uses over $1$M concept labels, whereas LICORICE achieves similar performance while using 2000×2000$\times$2000 × fewer labels.
Learning Concepts with Human Feedback. Prior research has explored leveraging human concept labels, but not for RL and without focusing on reducing the labeling burden. For instance, Lage & Doshi-Velez (2020) has users label additional information about the relevance of certain feature dimensions to the concept label to facilitate concept learning for high-dimensional data. In a different vein, Chauhan et al. (2023) develop a test-time intervention policy for querying for concept labels to improve the final prediction. These methods do not directly tackle the issue of reducing the overall labeling burden during training and could be used alongside our method. A concurrent work (Shao et al., 2024) also studies sample efficiency but mainly focus on human task demonstrations.
@@ 6 Discussion and Conclusion
In this work, we proposed LICORICE, a novel RL algorithm that addresses the critical issue of model interpretability while minimizing reliance on continuous human annotation. We introduced a training scheme that enables RL algorithms to learn concepts efficiently from little labeled concept data. We demonstrated how this approach reduces manual labeling effort. We conducted initial experiments to demonstrate how we can leverage powerful VLMs to infer concepts from raw visual inputs without explicit labels in some environments. There are broader societal impacts of our work that must be considered, including both the impacts of using VLMs in real-world applications and more general considerations around interpretability. For a more detailed discussion of limitations and future work, please refer to Appendix C.
Acknowledgments
This work was supported in part by NSF grant IIS-2046640 (CAREER). Co-author Fang is supported in part by a Sloan Research Fellowship. We thank NVIDIA for providing computing resources. We thank Zhicheng Zhang for some helpful discussions, Naveen Raman for pointing us to relevant literature, and Rex Chen and Zhiyu (Zoey) Chen for their constructive comments on a draft of this work.
@@ Appendix A Experimental Result Reproducibility
In this section, we provide descriptions of our concept definitions, prompts, and other experimental details toward the goal of reproducibility.
A.1 Definitions of Concepts
@@
PixelCartPole Cart Position Continuous $(-2.4,2.4)$ $0.01$\pm$0.0$$
@@
DoorKey Agent Position x Discrete 5 $0.36$\pm$0.08$$
@@
DynamicObstacles Agent Position x Discrete 3 $0.03$\pm$0.06$$
@@
Boxing Agent Position x at Frame 1 Discrete 160 $0.81$\pm$0.14$$
Agent Position y at Frame 1 Discrete 210 $0.93$\pm$0.05$$
Enemy Position x at Frame 1 Discrete 160 $0.78$\pm$0.13$$
Enemy Position y at Frame 1 Discrete 210 $0.9$\pm$0.07$$
Agent Position x at Frame 2 Discrete 160 $0.67$\pm$0.06$$
Agent Position y at Frame 2 Discrete 210 $0.88$\pm$0.04$$
Enemy Position x at Frame 2 Discrete 160 $0.71$\pm$0.05$$
Enemy Position y at Frame 2 Discrete 210 $0.83$\pm$0.08$$
Pong Ball Position x Discrete 176 $0.99$\pm$0.00$$
Ball Position y Discrete 88 $0.98$\pm$0.00$$
Ball Velocity x Discrete 176 $0.95$\pm$0.00$$
Ball Velocity y Discrete 176 $0.80$\pm$0.00$$
Paddle Velocity Discrete 176 $0.78$\pm$0.05$$
Paddle Acceleration Discrete 352 $0.79$\pm$0.04$$
Paddle Jerk Discrete 704 $0.83$\pm$0.04$$
Table 4: Concepts and their possible values for all environments. For discrete concepts, we report the number of categories. We also provide the mean GPT-4o labeling error for each single concept, averaged across 5 seeds and corresponding to the same evaluation protocols as the ones with the largest budgets in Table 7. We use MSELoss for continuous values and 1 - accuracy for discrete ones, and in Boxing, we regard one prediction correct if it is within distance 5 to the ground truth. The $\pm$ [value] part shows the standard deviation. The average errors over concepts are $0.24$\pm$0.06$$, $0.3$\pm$0.02$$, $0.13$\pm$0.03$$, $0.82$\pm$0.06$$, and $0.88\pm 0.02$ respectively for the environments.
Table 4 provides more details regarding the concepts used in each environment, categorizing them by their names, types, and value ranges. For the PixelCartPole environment, all concepts such as Cart Position, Cart Velocity, Pole Angle, and Pole Angular Velocity are continuous. In contrast, the DoorKey environment features discrete concepts like Agent Position (x and y), Key Position (x and y), and Door Open status, each with specific value ranges. Similarly, the DynamicObstacles environment lists discrete concepts, including Agent and Obstacle positions, with corresponding value ranges.
Precise concept definitions are crucial for enabling correct agent behavior. We visualize the start configurations for DoorKey in Figure 6 to illustrate the importance of well-defined concepts in reinforcement learning. DoorKey includes a variety of initial states, each with different positions of the agent, key, and door. This diversity highlights a challenge in concept-based RL: how to define concepts that are both specific enough to be meaningful and general enough to apply across all possible scenarios. An agent trained with poorly defined concepts might perform well in some configurations but fail to generalize to others, leading to suboptimal performance in new or unseen environments. As a result, concept engineering is its own separate but important problem.
Figure 6: Example start configurations of the DoorKey environment, along with the associated concepts for each configuration. This shows that concepts must be general enough to apply to many different configurations. As a result, the agent cannot memorize the exact position of the door and key.
A.2 LICORICE Details.
@@
Architecture. The network begins with a CNN-based feature extractor $f_{\theta}:\mathcal{X}\to\mathbb{R}^{d}$, comprising three convolutional layers that map input state $x\in\mathcal{X}$ to a $d$-dimensional feature vector $h=f_{\theta}(x)$. We then introduce a linear layer $g_{\phi}:\mathbb{R}^{d}\to\mathbb{R}^{k}$ for concept prediction ($g$), mapping extracted features to $k$ concept values. For regression tasks, each of the $k$ predicted concept is represented as a real value; for classification tasks, each concept maps to one categorical value derived from a classification head. The action prediction component ($f$) is implemented as an MLP with two fully connected layers, each containing 64 neurons with Tanh activation, taking only $c$ as input to produce action logits $a=\text{MLP}(c)$. Notably, we share the CNN feature extractor $f_{\theta}$ between policy and value functions, a decision informed by improved performance observed in preliminary experiments.
Value-Based Methods.
@@
Feature Extractor. If we use the actor-critic paradigm, we propose to share a feature extractor between the policy and value networks, shown in Figure 7. Intuitively, this choice can offer several advantages compared with using image or predicted concepts as input for both networks. Sharing a feature extractor enables both networks to benefit from a common, rich representation of the input data, reducing the number of parameters to be trained. More importantly, it balances the updates of the policy and value networks. In experiments, we observed that directly using the raw image as input for both networks complicated policy learning. Conversely, relying solely on predicted concepts for the value network may limit its accuracy in value estimation, particularly if the concepts do not capture all the nuances relevant to the value predictions.
@@ A.3 VLM Details
    Agent Direction: Specify the direction the red triangle is facing, which is the orientation of the vertex (pointy corner) of the isosceles triangle. Choose from ‘right’, ‘down’, ‘left’, or ‘up’. Clarify that this direction is independent of movement options.
@@
    • Status: Assess whether the door is open or closed (closed means the door is visible as a whole yellow cell, while open means the door disappears visually). Report as ‘true’ for open and ‘false’ for closed. Double-check the door’s appearance to confirm if it is open or closed.
@@
    Each direction’s feasibility should be reported as ‘true’ if clear and within the grid, and ‘false’ otherwise.
@@
    Agent Direction: Specify the direction the red triangle is facing, which is the orientation of the vertex (pointy corner) of the isosceles triangle. Choose from ‘right’, ‘down’, ‘left’, or ‘up’. Clarify that this direction is independent of movement options.
@@
    Each direction’s feasibility should be reported as ‘true’ if clear and within the grid, and ‘false’ otherwise.
[…]
  • The white player’s x and y coordinates
  • The black player’s x and y coordinates
Please carefully determine the following values and give concise answers one by one. Make sure to return an estimated value for each one, even if the task may look challenging.
Follow the reporting format:
  • frame 1 white x: estimated_value
  • frame 1 white y: estimated_value
  • frame 1 black x: estimated_value
  • frame 1 black y: estimated_value
  • frame 2 white x: estimated_value
  • frame 2 white y: estimated_value
  • frame 2 black x: estimated_value
  • frame 2 black y: estimated_value
[…]
Please determine the following values and give concise answers one by one. Make sure to return an estimated value for each one, even if the task may look challenging.
  • x coordinate of the ball relative to the paddle in the first frame
  • y coordinate of the ball in the first frame
  • x velocity of the ball calculated from the first and second frames
  • y velocity of the ball calculated from the first and second frames
  • velocity of the paddle calculated from the first and second frames
  • acceleration of the paddle calculated from the first three frames
  • jerk of the paddle calculated from the four frames
Follow the reporting format in your response.
  • x coordinate of the ball: estimated_value
  • y coordinate of the ball: estimated_value
  • x velocity of the ball: estimated_value
  • y velocity of the ball: estimated_value
  • velocity of the paddle: estimated_value
  • acceleration of the paddle: estimated_value
  • jerk of the paddle: estimated_value
@@ A.4 Experimental Details
Behavior learning hyperparameters. For the PPO hyperparameters, we set $4\cdot 10^{6}$ total timesteps for PixelCartPole and DoorKey, $10^{6}$ for DynamicObstacles, $1.5\cdot 10^{7}$ for Boxing, and $10^{7}$ for Pong. For PixelCartPole, DoorKey, and DynamicObstacles, we use $8$ vectorized environments, horizon $T=4096$, $10$ epochs for training, batch size of $512$, learning rate $3\cdot 10^{-4}$, entropy coefficient $0.01$, and value function coefficient $0.5$. For Boxing and Pong, we use $8$ vectorized environments, horizon $T=1024$, $4$ epochs for training, batch size of $256$, learning rate $3\cdot 10^{-4}$, entropy coefficient $0.01$, and value function coefficient $0.5$. For all other hyperparameters, we use the default values from Stable Baselines 3 Raffin et al. (2021).
Concept learning hyperparameters. For the concept training, we set $100$ epochs with Adam optimizer with the learning rate linearly decaying from $3\cdot 10^{-4}$ to 00 for each iteration in PixelCartPole, Boxing, and Pong. In DoorKey and DynamicObstacles, we use the same optimizer and initial learning rate, yet set $50$ epochs instead and set early stopping with threshold linearly increasing from $10$ to $20$, to incentivize the concept network not to overfit in earlier iterations. The batch size is $32$. We model concept learning for PixelCartPole as a regression problem (minimizing mean squared error). We model concept learning for DoorKey, DynamicObstacles, Boxing,and Pong as classification problems.
LICORICE-specific hyperparameters. For LICORICE, we set the ratio for active learning $\tau=10$, batch size to query labels in the active learning module $b=20$, and the number of ensemble models $N=5$ for the first three environments. For the complex environments Boxing and Pong, we choose the ratio for active learning $\tau=4$, batch size to query labels in the active learning module $b=B_{m}/5$ and number of ensemble models $N=5$ to make a balance between performance and speed. The default number of iterations chosen in our algorithm is $M=4$ for PixelCartPole, $M=2$ for DoorKey, DynamicObstacles, and $M=5$ for Boxing and Pong. The sample acceptance rate $p=0.02$ for PixelCartPoleand $p=0.1$ for the other four environments. The Random-Q baseline uses the same sample acceptance rate $p$ as LICORICE. In the complex environments Boxing and Pong, we use both the KL-divergence penalty and PPO loss at the end of the algorithm to improve optimization with $\beta=0.01$, as mentioned in Section 3.
Computational resources. We use NVIDIA A6000 and NVIDIA RTX 6000 Ada Generation. Each of our training program uses less than 3GB GPU memory. For Boxingand Pong, each run takes less than 20 hours to finish. For PixelCartPole and DoorKey, each run takes less than 9 hours to finish. For DynamicObstacles, each run takes less than 2 hours to finish.
Experimental replicability. For all experiments, we use 5 seeds [$123$, $456$, $789$, $1011$, $1213$] to train the models. We then evaluate on the environment with seed $42$ with 100 episodes and take the average of the reward across the episodes and the concept error across the observations within each episode.
@@ Appendix B Additional Results
In this section, we present both the numerical results as tables from the figures in the main paper, in addition to experimental results not reported in the main paper.
@@ B.1 Balancing Concept Performance and Environment Reward
PixelCartPole Sequential-Q 0.232 472±0.085 242plus-or-minus0.2324720.085242\$0.232\,472$\pm$\$0.085\,242\$0.232 472 ± 0.085 242 0.102 57±0.037 255plus-or-minus0.102570.037255\$0.102\,57$\pm$\$0.037\,255\$0.102 57 ± 0.037 255
@@
DoorKey Sequential-Q 0.462 566±0.091 999plus-or-minus0.4625660.091999\$0.462\,566$\pm$\$0.091\,999\$0.462 566 ± 0.091 999 0.462 327±0.045 955plus-or-minus0.4623270.045955\$0.462\,327$\pm$\$0.045\,955\$0.462 327 ± 0.045 955
@@
DynamicObstacles Sequential-Q 0.788 348±0.440 724plus-or-minus0.7883480.440724\$0.788\,348$\pm$\$0.440\,724\$0.788 348 ± 0.440 724 0.014 136±0.009 627plus-or-minus0.0141360.009627\$0.014\,136$\pm$\$0.009\,627\$0.014 136 ± 0.009 627
@@
Boxing Sequential-Q 0.865 427±0.195 112plus-or-minus0.8654270.195112\$0.865\,427$\pm$\$0.195\,112\$0.865 427 ± 0.195 112 0.318 058±0.260 921plus-or-minus0.3180580.260921\$0.318\,058$\pm$\$0.260\,921\$0.318 058 ± 0.260 921
Disagreement-Q 0.825 683±0.143 459plus-or-minus0.8256830.143459\$0.825\,683$\pm$\$0.143\,459\$0.825 683 ± 0.143 459 0.201 04±0.212 373plus-or-minus0.201040.212373\$0.201\,04$\pm$\$0.212\,373\$0.201 04 ± 0.212 373
Random-Q 0.957 467±0.039 969plus-or-minus0.9574670.039969\$0.957\,467$\pm$\$0.039\,969\$0.957 467 ± 0.039 969 0.142 976±0.099 212plus-or-minus0.1429760.099212\$0.142\,976$\pm$\$0.099\,212\$0.142 976 ± 0.099 212
LICORICE 1.004 997±0.017 713plus-or-minus1.0049970.017713\$1.004\,997$\pm$\$0.017\,713\$1.004 997 ± 0.017 713 0.049±0.010 967plus-or-minus0.0490.010967\$0.049$\pm$\$0.010\,967\$0.049 ± 0.010 967
LICORICE+GPT-4o 0.221 577±0.009 677plus-or-minus0.2215770.009677\$0.221\,577$\pm$\$0.009\,677\$0.221 577 ± 0.009 677 0.976 101±0.009 001plus-or-minus0.9761010.009001\$0.976\,101$\pm$\$0.009\,001\$0.976 101 ± 0.009 001
CPM 1.0±0.052 538plus-or-minus1.00.052538\$1.0$\pm$\$0.052\,538\$1.0 ± 0.052 538 0.000 109±0.000 103plus-or-minus0.0001090.000103\$0.000\,109$\pm$\$0.000\,103\$0.000 109 ± 0.000 103
Pong Sequential-Q 0.98±0.017 924plus-or-minus0.980.017924\$0.98$\pm$\$0.017\,924\$0.98 ± 0.017 924 0.536 369±0.034 322plus-or-minus0.5363690.034322\$0.536\,369$\pm$\$0.034\,322\$0.536 369 ± 0.034 322
Disagreement-Q 0.942 857±0.092 857plus-or-minus0.9428570.092857\$0.942\,857$\pm$\$0.092\,857\$0.942 857 ± 0.092 857 0.566 152±0.053 357plus-or-minus0.5661520.053357\$0.566\,152$\pm$\$0.053\,357\$0.566 152 ± 0.053 357
Random-Q 0.942 143±0.045 617plus-or-minus0.9421430.045617\$0.942\,143$\pm$\$0.045\,617\$0.942 143 ± 0.045 617 0.604 967±0.043 498plus-or-minus0.6049670.043498\$0.604\,967$\pm$\$0.043\,498\$0.604 967 ± 0.043 498
LICORICE 0.996 333±0.018 424plus-or-minus0.9963330.018424\$0.996\,333$\pm$\$0.018\,424\$0.996 333 ± 0.018 424 0.243 221±0.162 857plus-or-minus0.2432210.162857\$0.243\,221$\pm$\$0.162\,857\$0.243 221 ± 0.162 857
LICORICE+GPT-4o 0.126 19±0.170 042plus-or-minus0.126190.170042\$0.126\,19$\pm$\$0.170\,042\$0.126 19 ± 0.170 042 0.706 473±0.009 016plus-or-minus0.7064730.009016\$0.706\,473$\pm$\$0.009\,016\$0.706 473 ± 0.009 016
CPM 0.997 222±0.004 811plus-or-minus0.9972220.004811\$0.997\,222$\pm$\$0.004\,811\$0.997 222 ± 0.004 811 0.004 985±0.007 178plus-or-minus0.0049850.007178\$0.004\,985$\pm$\$0.007\,178\$0.004 985 ± 0.007 178
Table 5: Evaluation of the reward $R$ and concept error achieved by all methods in all environments. This is a extended table from Figure 2. The reward is reported as the fraction of the reward upper bound. For PixelCartPole, the $c$ error is the MSE. For the other environments, the $c$ error is 1 - accuracy. The first five algorithms are given a budget of $B=[500,300,300,3000,5000]$ for each environment, from top to bottom; CPM is given an unlimited budget (in practice, it uses 4M, 4M, 1M, 15M, 10M concept labels respectively). The $\pm$ [value] part shows the standard deviation. For each environment, we bold the highest reward and lowest concept error among the budget-constrained methods.
In Table 5, we present all of the numerical results from Figure 2, including standard deviation. Our method enjoys low variance across all environments in terms of both concept error and reward.